OpenStudy (dengeki_daisy):

I need help with this. A homeowner wants to build a circular patio. If the diameter of the patio is 20ft, what is its area to the nearest whole number. a)1,257 b)314 c)63 d)31

OpenStudy (mathmale):

Let me ask you: What's the area of a circle, if you're given the radius? What's the area, if you're given the diameter? Please type out the pertinent formulas. Thx.

OpenStudy (dengeki_daisy):

A = pi * r^2 A = 3.14 * 10^2 A = ?